Emmett:

A car travelling at a speed of 140 kmph covers 420 km. How much time will it take to cover 280 km?

Dva:

First, we need to find the time required to cover 420 km.

Dva:

Speed = Distance/Time

Emmett:

So 140 = 420/T and the t would equal 3 hours

Gucchi:

snow this is all you

Dva:

420 km = 3 hours so 280 km = (3/420) x 280 would equal what?

Emmett:

2

Emmett:

Would 2 be the answer?

AZ:

@emmett wrote:
Would 2 be the answer?
That's correct If they're driving at 140 kilometers per hour and they drive 420 km It would take them 3 hours to cover that distance 420 / 140 = 3 and so for them to cover 280 km at a speed of 140 kilometers per hour, 280 / 140 = 2 hours which is the answer you correctly arrived at